OpenStudy (wesdg1978):

Suppose that the radius r and the circumference C=2πr of a circle are differentiable functions of t. How do you write an equation that relates dC/dt to dr/dt?

OpenStudy (amistre64):

C = 2pi r C' = 2pi r'

OpenStudy (amistre64):

i think they have the setup\[\frac{dC}{dt}=\frac{dC}{dr}\frac{dr}{dt}\] but it all amounts to the same thing

OpenStudy (wesdg1978):

@amistre64 The available answers are: A) dr/dt = 2pi dC/dt B) dC/dt =dr/dt C) dC/dt = 2pir dr/dt D) dC/dt = 2pi(dr/dt)

OpenStudy (amistre64):

drop the fractional notation and stick with the prime notation ....

OpenStudy (amistre64):

C' = 2pi r' is in there